OpenStudy (anonymous):

what is the greatest common factor of the numbers using prime factorization inc. whether the numbers are relatively prime: 168, 180 and 450?

OpenStudy (anonymous):

clearly not relatively prime because they are all even

OpenStudy (anonymous):

\[168=2^3\times 3\times 7\] \[180=2^2\times 3^2\times 5\] \[450=2\times 3^2\times 5^2\]

OpenStudy (anonymous):

for least common multiple you will need \[2^3\] \[3^2\] \[5^2\] and \[7\] each factor you see to the highest power you see it in any ONE number

OpenStudy (anonymous):

so LCM is \[2^2\times 3^2\times 5^2\times 7\] whatever that is

OpenStudy (anonymous):

ok, thanks

OpenStudy (anonymous):

yw